# Satisficing Exploration in Bandit Optimization

### v1

@@
Satisficing Exploration in Bandit Optimization
@@
Satisficing Exploration in Bandit Optimization
@@
Motivated by the concept of satisficing in decision-making, we consider the problem of satisficing exploration in bandit optimization. In this setting, the learner aims at selecting satisficing arms (arms with mean reward exceeding a certain threshold value) as frequently as possible. The performance is measured by satisficing regret, which is the cumulative deficit of the chosen arm’s mean reward compared to the threshold. We propose SELECT, a general algorithmic template for Satisficing Exploration via LowEr Confidence bound Testing, that attains constant satisficing regret for a wide variety of bandit optimization problems in the realizable case (i.e., a satisficing arm exists). Specifically, given a class of bandit optimization problems and a corresponding learning oracle with sub-linear (standard) regret upper bound, SELECT iteratively makes use of the oracle to identify a potential satisficing arm with low regret. Then, it collects data samples from this arm, and continuously compares the LCB of the identified arm’s mean reward against the threshold value to determine if it is a satisficing arm. As a complement, SELECT also enjoys the same (standard) regret guarantee as the oracle in the non-realizable case. Finally, we conduct numerical experiments to validate the performance of SELECT for several popular bandit optimization settings.
@@ 1 Introduction
Multi-armed bandit (MAB) is a classic framework for decision-making under uncertainty. In MAB, a learner is given a set of arms with initially unknown mean rewards. She repeatedly picks an arm to collect its noisy reward and her goal is to maximize the cumulative rewards. The performance is usually measured by the notion of regret, which is the difference between the maximum possible total mean rewards and the total mean rewards collected by the learner. With its elegant form, MAB has found numerous applications in recommendation systems, ads display, and beyond, and many (nearly) optimal algorithms have been developed for various bandit optimization settings (Lattimore and Szepesvári 2020). However, as one can expect, in many MAB settings, especially those with complicated decision landscapes, major regret is unavoidable because finding the optimal action typically requires a lot of exploration.
@@
While making the optimal decision is desired in certain high-stakes situations, it is perhaps not surprising that what we often need is just a good enough or satisficing alternative (Lufkin 2021). Coined by Nobel laureate Herbert Simon (Simon 1956), the term “satisficing” refers to the decision-making strategy that settles for an acceptable solution rather than exhaustively hunting for the best possible. For example,
1) To avoid customer disengagement, retailers usually keep safety stock for every product to improve service level, i.e., the probability of no stockout (Jenkins 2022). Due to various practical constraints (e.g., warehouse capacity, budgets, demand variations), it can be challenging to compute the optimal combination of service levels. Therefore, retailers often choose to maintain a target service level for each of the products (see, e.g., Willemain 2017, Vandeput 2021);
2) In product design, there are always a lot of decisions to make, and it can take a long time if we try to optimize for every one of them. Hence, instead of finding the optimal design, it is more important to identify a design that meets the needs of a considerable portion of the targeted customers (see, e.g., Mabey 2022, Whitenton 2014);
3) In choosing a restaurant, a person is usually satisfied with one that meets certain criteria, such as decent food quality, reasonable price, and nearby locations, rather than finding the “best” restaurant that excels in all possible aspects (see, e.g., Aransiola 2023).
@@
Following the idea of satisficing, existing works have turned attention to satisficing bandits where the objective is to hit a satisficing arm (i.e., an arm whose mean reward exceeds a certain threshold value) as frequently as possible (Tamatsukuri and Takahashi 2019, Michel et al. 2023). The corresponding performance measure is satisficing regret, which is the cumulative deficit of the chosen arm’s mean reward compared to the threshold. In particular, Michel et al. (2023) develops SAT-UCB, an algorithm for finite-armed satisficing bandits: In the realizable case (i.e., a satisficing arm exists), the algorithm attains a constant regret, i.e., independent of the length of the time horizon; Otherwise, in the non-realizable case (e.g., we are in a high-stakes situation or the threshold value is out of reach), it preserves the optimal logarithmic (standard) regret.
@@
Main Contributions:
In this work, we propose SELECT, a novel algorithmic template for Satisficing Exploration via LowEr Confidence bound Testing. More specifically:
@@
  • In Section 4, we establish that in the realizable case or when a satisficing arm exists, SELECT is able to achieve a constant satisficing regret. Notably, with the design of forced sampling and LCB tests, the satisficing regret bound has no dependence on the satisficing gap (see the forthcoming Remark 2 for a detailed discussion). Instead, it scales inverse proportionally to the exceeding gap, which measures the difference between the optimal reward and the threshold value, and is positive in general; Otherwise (i.e., when no satisficing arm exists), SELECT enjoys the same standard regret bound as the learning oracle.
@@ 1.1 Related Works
Besides Michel et al. (2023), the most relevant works to ours are Bubeck et al. (2013), Garivier et al. (2019), Tamatsukuri and Takahashi (2019), all of which focus on finite-armed bandits. Bubeck et al. (2013), Garivier et al. (2019) introduce constant regret algorithms with knowledge of the optimal reward. This can be viewed as a special case of satisficing where the threshold value is exactly the optimal mean reward. Tamatsukuri and Takahashi (2019) also establishes constant regret for finite-armed satisficing bandits when there exists exactly one satisficing arm. As a follow-up paper of Michel et al. (2023), Hajiabolhassan and Ortner (2023) considers satisficing in finite-state MDP.
@@ 2 Problem Formulation
In this section, we present the setup of our problem. For a class of bandit optimization problems $\mathcal{P}$, the learner is given a set of feasible arms $\mathcal{X}$, selecting arm $X\in\mathcal{X}$ in a time step brings a random reward with mean $r(X)$, which is initially unknown to the learner. In each time step $t=1,2,\ldots,T$, the learner selects an arm $X_{t}\in\mathcal{X}$, and then observes and receives a noisy reward $Y_{t}=r(X_{t})+\epsilon_{t}$, where $\epsilon_{t}$ is a conditionally 1-subgaussian random variable, i.e., $\mathbb{E}[\epsilon_{t}|X_{1},\epsilon_{1}\ldots,X_{t-1},\epsilon_{t-1},X_{t}]=0$ and $\mathbb{E}[e^{\lambda\epsilon_{t}}|X_{1},\epsilon_{1}\ldots,X_{t-1},\epsilon_{t-1},X_{t}]\leq e^{\lambda^{2}/2}$ holds for all $\lambda\in\mathbb{R}$.
@@
We say that the setting is realizable if $r(X^{*})\geq S$; Otherwise, it is non-realizable. We recall that an arm $X$ is satisficing if $r(X)\geq S$; Otherwise, it is non-satisficing. Existing works on similar settings (e.g., Michel et al. (2023)) usually derive satisficing regret bound that scales inversely proportional to the satisficing gap
@@
Remark 1 (On the Definition of Satisficing Regret).
In some cases, the learner might only care about whether the total mean reward exceeds a certain threshold $S^{\prime}$. By the Jensen’s inequality and directly setting $S=S^{\prime}/T$, we can see that any upper bound on satisficing regret w.r.t. the threshold $S$ also implies a loss upper bound on total mean reward w.r.t. $S^{\prime}$.
3 SELECT: Satisficing Exploration via Lower Confidence Bound Testing
@@
When the arm space $\mathcal{X}$ is not too large, one can collect enough data samples for every arm to estimate its mean reward and gradually abandon any non-satisficing arms (see, e.g., Garivier et al. (2019), Michel et al. (2023)). In general, however, the arm space can be large and may even contain infinitely many arms (e.g., continuous arm space). Therefore, it becomes impossible to identify all non-satisficing arms. We thus take an alternative approach: It repeatedly locates a potentially satisficing arm with low (satisficing) regret. Then, it tests if this candidate arm is truly a satisficing arm or not. If not, it kicks off the next round of search.
To ensure fast detection of candidate arms with low regret for a given bandit problem class $\mathcal{P}$, SELECT makes use of a bandit algorithm for (standard) regret minimization on $\mathcal{P}$. To this end, we assume the access to a blackbox learning oracle that achieves sub-linear standard regret for the bandit problem class $\mathcal{P}$.
Assumption 1.
There exists a sequence of learning algorithms ALG such that for any given time horizon $t$, the regret of $\texttt{ALG}(t)$ under any instance of the problem class $\mathcal{P}$ is bounded by
@@
where $C_{1}\geq 1,1/2\leq\alpha<1,\beta\geq 0$ are constants only dependent on the problem class.
We remark that Assumption 1 only asks for sub-linearity in standard regret upper bounds rather than optimality. Therefore, it can be easily satisfied by a wide range of bandit optimization problems and the corresponding bandit learning algorithms.
With this, for a problem class $\mathcal{P}$ and its sub-linear regret learning algorithm ALG, SELECT runs in rounds/phases. In round $i$, it takes the following steps (see Fig. 1 for a illustration):
@@
Step 1. Identify a Potentially Satisficing Arm: Let $\gamma_{i}=2^{-i(1-\alpha)/\alpha}$, SELECT follows ALG for the first $t_{i}=\lceil\gamma_{i}^{-1/(1-\alpha)}\rceil=\lceil 2^{i/\alpha}\rceil$ steps of round $i$ and records its arm selections. Then, it samples an arm $\hat{X}_{i}$ uniformly random from the trajectory of ALG in this round. By virtue of ALG, we find an arm whose mean reward is at most $\tilde{O}(t^{\alpha-1}_{i})=\tilde{O}(\gamma_{i})$ below $r(X^{*})$ in expectation, and only $\tilde{O}(t^{\alpha}_{i})$ (satisficing) regret is incurred in this step. We note that in the realizable case, as $i$ increases, $\tilde{O}(\gamma_{i})$ will gradually become smaller than $\Delta_{S}^{*}$, meaning that the sampled arm $\hat{X}_{i}$ is more likely to be a satisficing arm and enjoys no satisficing regret;
Step 2. Forced Sampling on the Identified Arm: To validate if $\hat{X}_{i}$ is a satisficing arm or not, we need to collect enough data from pulling $\hat{X}_{i}$ to perform some statistical test on its mean reward. However, when the arm set $\mathcal{X}$ is large, ALG might only have pulled $\hat{X}_{i}$ for limited number of times. To circumvent this, we perform a forced sampling on $\hat{X}_{i}$ by pulling it for $T_{i}=\lceil\gamma^{-2}_{i}\rceil$ times;
Step 3. Lower Confidence Bound Tests: We are now ready to perform the statistical test. In this step, SELECT continues to collect data from $\hat{X}_{i}$. We use $k$ (initialized to 0) to denote the number of additional pieces of noisy reward data generated by pulling $\hat{X}_{i}$ in the current time step and $\hat{r}^{\text{tot}}_{i}$ to denote the running total reward collected from Step 2 and the current step so far. At the same time, it keeps comparing the LCB of $\hat{X}_{i}$’s mean reward, i.e., $\hat{r}^{\text{tot}}_{i}/(T_{i}+k)-\sqrt{4\log(T_{i}+k)/(T_{i}+k)}$, against the threshold value $S$ after acquiring every piece of new data. SELECT terminates this round and enters the next whenever the LCB is less than $S$.
@@
We now take a pause to comment on the design of Steps 2 and 3.
1) In Step 3, SELECT compares the LCB of $\hat{X_{i}}$’s mean reward against $S$ to determine if $X_{i}$ is a satisficing arm. This deviates from prior works that use UCB Garivier et al. (2019) or empirical mean Michel et al. (2023). However, it turns out that this is essential in achieving a constant satisficing regret that does not scale with $1/\Delta_{S}$, which can quickly explode in many cases. As we will see, once entering Step 3, SELECT can terminate a round within 1 additional time step in expectation when facing a non-satisficing arm (i.e., (3) in the proof of the forthcoming Proposition 1). In contrast, if it follows prior works to use UCB or empirical mean instead, the number of time steps required (and hence, the satisficing regret) will unavoidably scale with $1/\Delta_{S}$ (see e.g., Theorem 9 of Garivier et al. (2019) or Theorem 1 of Michel et al. (2023));
2) One challenge from using the LCB test in Step 3 is that it is a more conservative design compared to prior works. In fact, if we directly enter Step 3 without Step 2, the LCB of $r(\hat{X_{i}})$ can easily fall below $S$ even if $r(\hat{X}_{i})\geq S$. This is because we may not have pulled $\hat{X}_{i}$ for enough times in Step 1 and the corresponding confidence interval might be large. As a result, the value of $r(\hat{X}_{i})$ LCB, which is the difference between empirical mean and confidence interval, can be extremely small. This indicates, without the forced sampling in Step 2, major satisficing regret can be incurred due to frequent re-start of a new round. With the forced sampled data collected from Step 2, SELECT ensures that the width of the confidence interval is of order $\tilde{O}(T_{i}^{-1/2})=\tilde{O}(\gamma_{i})$ (i.e., same as $\mathbb{E}[r(X^{*})-r(\hat{X}_{i})]$). Consequently, whenever $\gamma_{i}$ shrinks to well below $\Delta_{S}^{*}$, SELECT gradually becomes less likely to terminate a round (see the forthcoming Proposition 2).
Algorithm 1 Satisficing Exploration for General Bandit (SELECT)
[…]

### v2

# Satisficing Regret Minimization in Bandits

@@
Satisficing Regret Minimization in Bandits
@@
Satisficing Regret Minimization in Bandits
@@
Motivated by the concept of satisficing in decision-making, we consider the problem of satisficing regret minimization in bandit optimization. In this setting, the learner aims at selecting satisficing arms (arms with mean reward exceeding a certain threshold value) as frequently as possible. The performance is measured by satisficing regret, which is the cumulative deficit of the chosen arm’s mean reward compared to the threshold. We propose SELECT, a general algorithmic template for Satisficing REgret Minimization via SampLing and LowEr Confidence bound Testing, that attains constant satisficing regret for a wide variety of bandit optimization problems in the realizable case (i.e., a satisficing arm exists). Specifically, given a class of bandit optimization problems and a corresponding learning oracle with sub-linear (standard) regret upper bound, SELECT iteratively makes use of the oracle to identify a potential satisficing arm with low regret. Then, it collects data samples from this arm, and continuously compares the LCB of the identified arm’s mean reward against the threshold value to determine if it is a satisficing arm. As a complement, SELECT also enjoys the same (standard) regret guarantee as the oracle in the non-realizable case. Finally, we conduct numerical experiments to validate the performance of SELECT for several popular bandit optimization settings.
@@ 1 Introduction
Multi-armed bandit (MAB) is a classic framework for decision-making under uncertainty. In MAB, a learner is given a set of arms with initially unknown mean rewards. She repeatedly picks an arm to collect its noisy reward and her goal is to maximize the cumulative rewards. The performance is usually measured by the notion of regret, which is the difference between the maximum possible total mean rewards and the total mean rewards collected by the learner. With its elegant form, MAB has found numerous applications in recommendation systems, ads display, and beyond, and many (nearly) optimal algorithms have been developed for various bandit optimization settings (Lattimore and Szepesvári 2020).
@@
While making the optimal decision is desired in certain high-stakes situations, it is perhaps not surprising that what we often need is just a good enough or satisficing alternative (Lufkin 2021). Coined by Nobel laureate Herbert Simon (Simon 1956), the term “satisficing” refers to the decision-making strategy that settles for an acceptable solution rather than exhaustively hunting for the best possible.
The concept of satisficing finds its place in many real world decision-making under uncertainty settings (Caplin et al. 2011). For example, in inventory management, studies show that product stockouts can result in customer disengagement and impair a firm’s profitability (Fitzsimons 2000, Jing and Lewis 2011). As such, retailers usually replenish their products periodically to improve service level (i.e., probability of no stockout for any product). Notably, due to various constraints (e.g., capacity and inventory cost), instead of blindly maximizing service level, retailers typically set a target service level (e.g., the ideal service level in retailing is $90\%\sim 95\%$ (Levy et al. 2023)). Consequently, in face of the unknown demand distributions, one can cast the problem into a MAB setting with a different objective. In this setting, a retailer starts without complete knowledge of the demand distributions and proceeds by trying various stock level combinations (i.e., arms) of the products over time. Here, the objective is to hit the target service level as frequently as possible rather than maximizing the service level every time. As another example, in fashion product design, designers usually need to design a sequence of products for a group of customers with initially unknown preferences. Instead of demanding for the best design every time, designers only need to pick a design that can hopefully meet the expectation of a certain portion of the customers for most of the time (Mabey 2022, Whitenton 2014).
@@
The above examples give rise to the study of satisficing bandits (Tamatsukuri and Takahashi 2019, Michel et al. 2023) where the objective is to hit a satisficing arm (i.e., an arm whose mean reward exceeds a certain threshold value) as frequently as possible. The corresponding performance measure is satisficing regret, which is the cumulative deficit of the chosen arm’s mean reward compared to the threshold. In particular, Michel et al. (2023) develops SAT-UCB, an algorithm for finite-armed satisficing bandits: In the realizable case (i.e., a satisficing arm exists), the algorithm attains a constant satisficing regret, i.e., independent of the length of the time horizon; Otherwise, in the non-realizable case (i.e., the threshold value is out of reach), it preserves the optimal logarithmic (standard) regret.
@@
Main Contributions: In this work, we propose SELECT, a novel algorithmic template for Satisficing REgret Minimization via SampLing and LowEr Confidence Bound Testing. More specifically:
@@
  • In Section 4, we establish that in the realizable case or when a satisficing arm exists, SELECT is able to achieve a constant satisficing regret. Notably, the satisficing regret bound has no dependence on the satisficing gap (see the forthcoming Remark 2 for a detailed discussion). Instead, it scales inverse proportionally to the exceeding gap, which measures the difference between the optimal reward and the threshold value, and is positive in general; Otherwise (i.e., when no satisficing arm exists), SELECT enjoys the same standard regret bound as the learning oracle.
@@ 1.1 Related Works
Besides Michel et al. (2023), the most relevant works to ours are Bubeck et al. (2013), Garivier et al. (2019), Tamatsukuri and Takahashi (2019), all of which focus on finite-armed bandits. Bubeck et al. (2013) and Garivier et al. (2019) introduce constant regret algorithms with knowledge of the optimal reward. This can be viewed as a special case of satisficing where the threshold value is exactly the optimal mean reward. Hüyük and Tekin (2021) further extend the algorithm in Garivier et al. (2019) to multi-objective multi-armed bandit settings. Tamatsukuri and Takahashi (2019) also establishes constant regret for finite-armed satisficing bandits when there exists exactly one satisficing arm. As a follow-up paper of Michel et al. (2023), Hajiabolhassan and Ortner (2023) considers satisficing in finite-state MDP.
@@
The notion of satisficing is also studied in the pure exploration setting (Locatelli et al. 2016, Mukherjee et al. 2017, Kano et al. 2019) as thresholding bandits. Their objective is to identify all arms with mean reward above a certain satisficing level with limited exploration budget. Unlike the satisficing regret we consider, the performance of their algorithm is measured by simple regret, i.e., the expected number of misidentification made by the final output. Note that any pure exploration algorithm (Audibert and Bubeck 2010) is able to identify the optimal arm or a satisficing arm with high probability after a number of steps. However, due to a small but positive error probability, subsequent exploitation of the identified arm will always incur linear satisficing regret, thus a simple explore-then-exploit approach is unable to obtain a constant satisficing regret bound. Similar observations have also been made in Michel et al. (2023).
@@ 2 Problem Formulation
In this section, we present the setup of our problem. Consider a class of bandit optimization problem $(\mathcal{X},\mathcal{R})$, where $\mathcal{X}$ is the arm set and $\mathcal{R}\subseteq\{r:\mathcal{X}\to\mathbb{R}\}$ is the class of admissible reward functions. The learner is given a set of feasible arms $\mathcal{X}$ and a class of admissible reward functions $\mathcal{R}$, pulling arm $X\in\mathcal{X}$ in a time step brings a random reward with mean $r(X)$. The learner knows the underlying reward function $r(\cdot)$ belongs to the function class $\mathcal{R}$, but the exact $r(\cdot)$ is initially unknown to the learner.
In each time step $t=1,2,\ldots,T$, the learner pulls an arm $X_{t}\in\mathcal{X}$, and then observes and receives a noisy reward $Y_{t}=r(X_{t})+\epsilon_{t}$, where $\epsilon_{t}$ is a conditionally 1-subgaussian random variable, i.e., $\mathbb{E}[\epsilon_{t}|X_{1},\epsilon_{1}\ldots,X_{t-1},\epsilon_{t-1},X_{t}]=0$ and $\mathbb{E}[e^{\lambda\epsilon_{t}}|X_{1},\epsilon_{1}\ldots,X_{t-1},\epsilon_{t-1},X_{t}]\leq e^{\lambda^{2}/2}$ holds for all $\lambda\in\mathbb{R}$.
Remark 1.
[…]
@@
We say that the setting is realizable if $r(X^{*})\geq S$; Otherwise, it is non-realizable. We recall that an arm $X$ is satisficing if $r(X)\geq S$; Otherwise, it is non-satisficing. Existing works on similar settings (e.g., Michel et al. 2023) usually derive satisficing regret bound that scales inversely proportional to the satisficing gap
@@
3 SELECT: Satisficing Regret Minimization via Sampling and Lower Confidence Bound Testing
@@
When the arm space $\mathcal{X}$ is not too large, one can collect enough data samples for every arm to estimate its mean reward and gradually abandon any non-satisficing arms (see, e.g., Garivier et al. 2019, Michel et al. 2023). In general, however, the arm space can be large and may even contain infinitely many arms (e.g., continuous arm space). Therefore, it becomes impossible to identify all non-satisficing arms. We thus take an alternative approach: It repeatedly locates a potentially satisficing arm with low (satisficing) regret. Then, it tests if this candidate arm is truly a satisficing arm or not. If not, it kicks off the next round of search.
To ensure fast detection of candidate arms with low regret for a given bandit problem class $(\mathcal{X},\mathcal{R})$, SELECT makes use of a bandit algorithm for (standard) regret minimization on $(\mathcal{X},\mathcal{R})$. To this end, we assume access to a blackbox learning oracle that achieves sub-linear standard regret for the bandit problem class $(\mathcal{X},\mathcal{R})$.
Condition 1.
For problem class $(\mathcal{X},\mathcal{R})$, there exists a sequence of learning algorithms ALG such that for any reward function $r\in\mathcal{R}$ and any time horizon $t\geq 2$, the regret of $\texttt{ALG}(t)$ is bounded by
@@
where $C_{1}\geq 1,1/2\leq\alpha<1,\beta\geq 0$ are constants only dependent on the problem class and independent from the time horizon $t$.
We remark that Condition 1 only asks for sub-linearity in standard regret upper bounds rather than optimality. Therefore, it is an extremely mild condition and can be easily satisfied by a wide range of bandit optimization problems and the corresponding bandit learning oracles, e.g., finite-armed bandits, Lipschitz bandits, etc. We will demonstrate this in the forthcoming Section 5 and Appendix D.
With this, for a problem class $(\mathcal{X},\mathcal{R})$ and its sub-linear regret learning algorithm ALG, SELECT runs in rounds. In round $i$, it takes the following steps (see Fig. 1 for a illustration):
@@
Step 1. Identify a Potentially Satisficing Arm: Let $\gamma_{i}=2^{-i(1-\alpha)/\alpha}$, SELECT follows $\texttt{ALG}(t_{i})$ for the first $t_{i}=\lceil\gamma_{i}^{-1/(1-\alpha)}\rceil=\lceil 2^{i/\alpha}\rceil$ steps of round $i$ and records its arm selections. Then, it samples an arm $\hat{X}_{i}$ uniformly random from the trajectory of $\texttt{ALG}(t_{i})$ in this round. By virtue of ALG, we find an arm whose mean reward is at most $\tilde{O}(t^{\alpha-1}_{i})=\tilde{O}(\gamma_{i})$ below $r(X^{*})$ in expectation, and only $\tilde{O}(t^{\alpha}_{i})$ satisficing regret is incurred in this step in the realizable case. We note that in the realizable case, as $i$ increases, $\tilde{O}(\gamma_{i})$ will gradually become smaller than $\Delta_{S}^{*}$, meaning that the sampled arm $\hat{X}_{i}$ is more likely to be a satisficing arm and enjoys no satisficing regret;
Step 2. Forced Sampling on the Identified Arm: To validate if $\hat{X}_{i}$ is a satisficing arm or not, we need to collect enough data from pulling $\hat{X}_{i}$ before performing any statistical test (see the forthcoming Remark 2 for the reasons). However, when the arm set $\mathcal{X}$ is large, $\texttt{ALG}(t_{i})$ might only have pulled $\hat{X}_{i}$ for limited number of times. To circumvent this, we perform a forced sampling on $\hat{X}_{i}$ by pulling it for $T_{i}=\lceil\gamma^{-2}_{i}\rceil$ times;
Step 3. Lower Confidence Bound Tests: In this step, SELECT kicks off the LCB test as more data is collected from pulling $\hat{X}_{i}$. We use $k$ (initialized to 0) to denote the number of additional pieces of noisy reward data generated by pulling $\hat{X}_{i}$ in the current time step and $\hat{r}^{\text{tot}}_{i}$ to denote the running total reward collected from Step 2 and the current step so far. At the same time, it keeps comparing the LCB of $\hat{X}_{i}$’s mean reward, i.e., $\hat{r}^{\text{tot}}_{i}/(T_{i}+k)-\sqrt{4\log(T_{i}+k)/(T_{i}+k)}$, against the threshold value $S$ after acquiring every piece of new data. SELECT terminates this round and enters the next whenever the LCB is less than $S$.
@@
(Key Novelties of SELECT) Our algorithm brings together several key ingredients, including sampling from oracle trajectories (Step 1), forced sampling to collect data (Step 2), and the LCB test (Step 3). In what follows, we further comment on the subtleties of each step.
Step 1. Prior works on satisficing (Garivier et al. 2019, Michel et al. 2023) mainly use uniform exploration to find candidates of satisficing arms. This approach could incur major satisficing regret when the arm set is large or even possibly infinite, as conducting uniform exploration on a large arm set may result in pulling a large number of non-satisficing arms. To bypass this challenge, in Step 1, we search for candidates using a bandit learning oracle with sub-linear standard regret. Hence, we are able to find candidates of satisficing arms without incurring large regret.
Step 2. One challenge from using the LCB test in Step 3 is that it is a more conservative design compared to prior works. In fact, if we directly enter Step 3 without Step 2, the LCB of $r(\hat{X_{i}})$ can easily fall below $S$ even if $r(\hat{X}_{i})\geq S$. This is because we may not have pulled $\hat{X}_{i}$ for enough times in Step 1 and the corresponding confidence interval might be large. As a result, the lower confidence bound of $r(\hat{X}_{i})$, which is the difference between empirical mean and confidence radius, can be significantly below $S$. This indicates, without the forced sampling in Step 2, major satisficing regret can be incurred due to frequent re-start of a new round. With the forced sampled data collected from Step 2, SELECT ensures that the width of the confidence interval is of order $\tilde{O}(T_{i}^{-1/2})=\tilde{O}(\gamma_{i})$ (i.e., same as $\mathbb{E}[r(X^{*})-r(\hat{X}_{i})]$). Consequently, whenever $\gamma_{i}$ shrinks to well below $\Delta_{S}^{*}$, SELECT gradually becomes less likely to terminate a round (see the forthcoming Proposition 2);
Step 3. In Step 3, SELECT compares the LCB of $\hat{X_{i}}$’s mean reward against $S$ to determine if $X_{i}$ is a satisficing arm. This deviates from prior works that use UCB (Garivier et al. 2019) or empirical mean (Michel et al. 2023). However, it turns out that this is essential in achieving a constant satisficing regret that does not scale with $1/\Delta_{S}$, which can quickly explode in many cases. As we will see, once entering Step 3, SELECT can terminate a round within 1 additional time step in expectation when facing a non-satisficing arm (i.e., (3) in the proof of the forthcoming Proposition 1). In contrast, if it follows prior works to use UCB or empirical mean instead, the number of time steps required (and hence, the satisficing regret) will unavoidably scale with $1/\Delta_{S}$ (see e.g., Theorem 9 of Garivier et al. 2019 or Theorem 1 of Michel et al. 2023).
Algorithm 1 Satisficing Regret Minimization via Sampling and Lower Confidence Bound Testing (SELECT)